# Robust Robotic Exploration and Mapping Using Generative Occupancy Map Synthesis

### v1

@@ 8.1.2 Narrow Hallways and Corners
In this scenario we can directly measure the impact that SceneSense has on robot navigation. We provide 2 experiments to evaluate the SceneSense occupancy enhancements when traversing narrow corners. First we run 10 corner traversals in the same location to evaluate consonant of the enhancements. Second, We select 10 different narrow hallway scenarios to evaluate the SceneSense predictions in different environments and generalization abilities. In both scenarios we set a start and end point during the hallway exploration and set the robot to explore. We evaluate the following systems for comparison.
@@
    GBPlanner [18]: Baseline GBPlanner is used to explore the narrow hallway environments.
@@
    Method Exp. Time Min Time Max Time $\boldsymbol{\sigma_{\text{exp}}}$ Failures
@@ 8.1.3 Discussion
Figure 11: Test environments for exploration evaluation. (a) Environment 1. (b) Environment 2. The green voxels represent the complete occupancy map, and the black cubes indicate the robot’s starting positions.
8.1.4 SceneSense Robustness
While promising, the results presented in Sections 8.1.1 and 8.1.2 could be solved with simple brute force solution such as removing the ground checking from the graph sampler. To further differentiate SceneSense from potential brute force solutions we identify adversarial scenarios where areas of the map are not traversable. One such scenario is shown in figure 8, where the robot traverses near a glass railing unobservable to the lidar. While a naive unconstrained sampling approach could sample points and paths above the empty space, SceneSense is able to differentiate this space from other potential ground areas and not predict ground over the edge.
@@ 8.2.1 Experimental Setup
Figure 12: (a) The vision-only map. (b) The SceneSense-enhanced map. The current planning graph is shown as the purple and red graph structures. In (a), the vision-only map contains a large gap near the robot’s starting position. Due to occlusions from the narrow hallway, the hole is not observed and therefore remains unpopulated. SceneSense, however, merges its predictions with the existing map to make a realistic estimation of the unobserved space, filling the holes and allowing the robot to plan over the environment. Traversable occlusions, such as vertical structures, are correctly negotiated out of the map when the robot makes observations of the space.
@@
To evaluate the exploration speed of the system we deployed the platform in in a test environments shown in Figure 11(a)and 11(b). The robot is started once at each position indicated by the black boxes in the associated figures (5 positions in environment 1, 4 positions in environment 2). One run per starting position is then run per configuration and the results are averaged to generate Figure 13 and Tables 3 and 4. The same testing configurations described in Section 8.1.2 are used for all runs.
@@
[…]
@@
[…]
@@ 8.2.2 Discussion
In both testing environments, the map enhanced with SceneSense outperforms the vision only map in average exploration time, max exploration time, and average exploration gain per second. Interestingly, the minimum exploration time is faster when operating over the vision only map in both cases, however the max time is also much larger. These results suggests lower consistency of operation when operating over vision only maps. In some starting configurations the robot may make observations that allow for rapid exploration, however in others the robot may make poor observations leading to increased total exploration time. Local occupancy prediction seems to narrow this distribution, resulting in a system that is less reliant on an optimal direct observation at startup.
Notably, frontier-centric SceneSense fails to complete the exploration of the space. While the initial rates of exploration with frontier-centric SceneSense are promising, the planning stacks inability to disambiguate predicted space and observed space leads to the planner believing the space is fully observed before exploration is actually complete. These results indicate a need for the development of new planners, that are able to distinguish observed space from predicted space and make coherent exploration decisions based on this additional information.
@@
The minimum time in environment 1 is attributed to the vision only GBPlanner implementation navigating straight down the hallway from the when starting from the dead-end instead of turning right at the first junction as shown in Figure 11(a). With the more complete ground representation, the SceneSense enhanced GBPlanner navigates down the larger hallway to the right first, increasing its overall completion time. Without this outlier existing in the environment 2 data collection, the exploration rate of the SceneSense enhanced planner is increased by $32\%$.
It is our hypothesis that the substantial exploration rate gains are actually attributed to the startup filling. As shown in Figure 12(b) the startup filling has a lingering effect on the map even after some initial exploration. It is likely the exploration gain in these areas is still significant, resulting in the planner choosing to double back to explore the starting position in several instances. Further, the existence of holes in the map can result in unintuitive planning results. As shown in Figure 14, a hole in the map leads to a path with a loop back around the hole, configuring the robot in a poor position to continue exploration. By filling these holes during occupancy prediction the planner has a cleaner space to generate potential exploration samples, as well as a reduction in overall explorable space in areas that have already been traversed.
@@ 9 Conclusion
20202020404040406060606080808080100100100100101010102020202030303030Denoising StepsFIDKID (×$\times 1000$)
22224444666688881010101010101010151515152020202025252525Guidance ScaleFIDKID (×$\times 1000$)
Cond = 1 Inpaint = 0 Cond = 0 Inpaint = 1 Cond = 1 Inpaint = 1 00202020204040404060606060Values
@@
This ablation study was first presented in [13] where the SceneSense occupancy prediction network was a conditional diffusion model (as opposed to the unconditional model presented here). When examining these results it was concluded that in this particular application the conditional network does not provided substantial benefits over the equivalent unconditional network.
@@
The number of diffusion steps defines the size of each diffusion step during the reverse diffusion process. Generating reasonable results using the fewest possible denoising steps is desirable behavior to reduce computation time. Additionally too many denoising steps have been shown to introduce sampling drift which results in decreased performance [44, 45]. As shown in 15 (a) our method saw the best results when configured to 30 denoising steps. Too few denoising steps results in the network being unable make accurate predictions over the large time step. Increasing the number of denoising steps keeps results relatively stable over time, however you can see the KID is slightly worse using more steps due to sampling drift. Sampling drift is a result of the discrepancy between the distribution of the training and the inference data. During training, the model is trained to reduce a noisy map $x_{t}$ to a ground truth map $x$, at inference time the model iteratively removes noise from its already imperfect noise predictions. These predictions will drift away from the initial corruption distribution which becomes more pronounced at smaller time steps due to the compounding error.
@@
The guidance scale $s$ as defined in [31] is a constant that multiplies the difference of the conditional diffusion and unconditional diffusion to “push” the diffusion process towards the conditioned answer. Setting $s$ too high results in too large of pushes away from reasonable predictions and results in poor generalization to new environments. The best FID is measured when $s=3$ however KID is slightly lower when $s=1$. Further, when examining chart (c) it is shown that the results when conditioning is removed all together ($s=0$ ) are very similar to the best results seen with conditioning enabled (albeit slightly worse). This is likely because most of the useful conditioning information is captured in the local occupancy data, and mapping measured RGB-D points from areas in front of the geometry to under or behind the platform is a very difficult task. The performance of the conditioning only data may be seen to have a larger impact on the overall results if the sensor could capture more local information, such as given a wider FOV or different mounting angle.
@@
This ablation study was first presented in [14] to evaluate the inference time gains of removing cross-attention based conditioning from the SceneSense occupancy prediction network.
In this ablation we implement the same denoising network structure presented in [13]. It is a U-net constructed from the HuggingFace Diffusers library of blocks [41] and consists of Resnet [42] downsampling/upsampling blocks. The diffusion model is trained using randomly shuffled pairs of ground truth local occupancy maps $x$. We use Chameleon cloud computing resources [48] to train our model on one A100 with a batch size of $32$ for $250$ epochs or $88,208$ training steps. We use a cosine learning rate scheduler with a 500 step warm up from $10^{-6}$ to $10^{-4}$. The noise scheduler for diffusion is set to $1,000$ noise steps.
@@
We evaluate the inference time of the unconditional diffusion model against the inference time of the conditional model presented in the original SceneSense paper [13]. The cross-attention enabling trainable parameters are removed for the unconditional model, but the number of output channels for the constructed U-net are held constant between both models. As the ablation results of the original paper show minor, or no performance gain between the conditional and unconditional model in this configuration we do not evaluate the results of the model predictions in these experiments.
@@
Cond. Model [13] Uncond. Model

### v2

@@ 8.1.2 Narrow Hallways and Corners
In this scenario we can directly measure the impact that SceneSense has on robot navigation. We provide 2 experiments to evaluate the SceneSense occupancy enhancements when traversing narrow corners. First, we run 10 corner traversals in the same location to evaluate consonant of the enhancements. Second, we select 10 different narrow hallway scenarios to evaluate the SceneSense predictions in different environments and generalization abilities. In both scenarios we set a start and end point during the hallway exploration and set the robot to explore. We evaluate the following systems for comparison.
@@
    GBPlanner [dang2020graph]: Baseline GBPlanner is used to explore the narrow hallway environments.
@@
    Method Exp. Time (s) Min Time (s) Max Time (s) $\boldsymbol{\sigma_{\text{exp}}}$ (s) Fail Count
@@ 8.1.3 Discussion
Figure 11: Two test environments for exploration evaluation. The green voxels represent the complete occupancy map, and the black cubes indicate the robot’s starting positions. (a) Environment 1 (b) Environment 2.
8.1.4 SceneSense Robustness to Adversarial Scenarios
While promising, the results presented in Sections 8.1.1 and 8.1.2 could be solved with simple brute force solution such as removing the ground checking from the graph sampler. To further differentiate SceneSense from potential brute force solutions we identify adversarial scenarios where areas of the map are not traversable. One such scenario is shown in Figure 7, where the robot traverses near a glass railing unobservable by the lidar sensor. While a naive unconstrained sampling approach could sample points and paths above the empty space, SceneSense is able to differentiate this space from other potential ground areas and not predict ground over the edge.
@@ 8.2.1 Experimental Setup
Figure 12: (a) The vision only map. (b) The SceneSense enhanced map. The current planning graph is shown as the purple and red graph structures. In (a), the vision only map contains a large gap near the robot’s starting position. Due to occlusions from the narrow hallway, the hole is not observed and therefore remains unpopulated. SceneSense, however, merges its predictions with the existing map to make a realistic estimation of the unobserved space, filling the holes and allowing the robot to plan over the environment. Traversable occlusions, such as vertical structures, are correctly negotiated out of the map when the robot makes observations of the space.
@@
To evaluate the exploration speed of the system we deployed the platform in test environments shown in Figures 11(a)and 11(b). The robot is started once at each position indicated by the black boxes in the associated figures (5 positions in environment 1, 4 positions in environment 2). One run per starting position is then run per configuration, and the results are averaged to generate Figure 13 and Tables 3 and 4. The same testing configurations described in Section 8.1.2 are used for all runs.
@@
Method Exp. Time (s) Min Time (s) Max Time (s) $\boldsymbol{\sigma_{\text{exp}}}$ (s) $\boldsymbol{\frac{\text{exp}\%}{\text{sec}}}$
@@
Method Exp. Time Min Time Max Time $\boldsymbol{\sigma_{\text{exp}}}$ $\boldsymbol{\frac{\text{exp}\%}{\text{sec}}}$
@@ 8.2.2 Discussion
Our experiments demonstrate that SceneSense-enhanced maps consistently improve exploration performance compared to vision only maps. In both test environments, SceneSense yields lower average and maximum exploration times with higher exploration gain per second. The vision only maps showed faster minimum exploration times, but with much larger maximum exploration times, indicating far greater variance. These results suggest lower consistency of operation when relying solely on vision. In some starting configurations the robot may make observations that enable rapid exploration, while in others it may make poor observations that increase total exploration time. Local occupancy prediction narrows this distribution, resulting in a system that is less dependent on a favorable direct observation at startup.
@@
Our data revealed that startup configurations lead to some environment specific behavior. The minimum time in environment 1 is attributed to the vision-only GBPlanner navigating straight down the hallway when starting from the dead end, instead of turning right at the first junction as seen in Figure 11(a). With the more complete ground representation from the SceneSense enhanced map, GBPlanner instead navigates down the larger hallway to the right first, which increases its overall completion time. Without this outlier in the environment 2 data collection, the exploration rate of the SceneSense enhanced planner increases by 32%. It is our hypothesis that the substantial exploration rate gains are actually attributed to the startup filling. As shown in Figure 12(b), the startup filling has a lingering effect on the map even after some initial exploration. It is likely that the exploration gain in these areas remains significant, resulting in the planner choosing to double back to explore the starting position in several instances.
We observed a clear distinction between local SceneSense predictions and frontier-centric SceneSense predictions. Local predictions improved planning performance both by filling in startup regions and by eliminating small holes in the map that otherwise can cause unintuitive behavior. For instance, in Figure 14, a small hole causes the planner to loop back on itself rather than taking a direct path. SceneSense fills such gaps, producing maps that enable coherent trajectory generation and thereby contributing to faster and more consistent exploration.
In contrast, frontier-centric SceneSense (SS frontier), while initially promising, was unable to complete exploration as seen in Figure 13. This failure arises because diffusion predictions overwrite unobserved voxels with occupancy probability values, which erases the boundary between free and unknown that frontier planners rely on. As a result, the planner prematurely concludes that the environment is fully explored. One possible mitigation is to add an indicator variable that marks whether a voxel was generated by diffusion or observed directly. While this allows for differentiation between diffused predicted voxels and sensor observed voxels, it does not recover the geometry of the frontier and would require the planner to inherently understand when to reason over flag based voxels and when to reason over the traditional sensor observed voxel representations. In practice, the planner would need to reason explicitly over the interface between diffusion-generated occupancies and those derived from sensor measurements in order to continue exploration using frontier based exploration. These results suggest that new planners are needed which can reason jointly about observed and predicted occupancy, preserving the structural cues necessary for frontier-based exploration.
Although SS frontier did not complete exploration, we argue it remains a valid and motivating contribution. First, it demonstrates the potential of extending SceneSense beyond local completions toward map-wide inference. Second, its failure case highlights the broader need for exploration planning paradigms that explicitly incorporate predictive models with joint reasoning, rather than treating predictions as direct drop in substitutions for perception. These insights point toward important research questions on integrating predicted occupancy into higher level reasoning and planning systems, which pursue in future work.
@@
9 Limitations
Despite the promising results demonstrated by our proposed method for generative occupancy map synthesis, several limitations warrant discussion and present opportunities for future work.
Currently, training and evaluation of SceneSense (SS) is limited to structured indoor environments. The generalization of our method to more diverse, complex scenarios, such as outdoor, unstructured, or highly dynamic environments, remains an open question. The training data, collected from a single building, may not fully capture the variability present in real-world conditions. Future work could train the model on broader datasets to improve its robustness and generalizability.
Our system’s reliance on an off-board GPU for inference presents a practical limitation for autonomous systems with strict power and size constraints. Deployment on embedded platforms would require model optimization and potentially an exploration of more efficient architectures.
Incorrect predictions, particularly in scenarios involving transparent or reflective surfaces, could produce unpredictable behavior. While the probabilistic map merging (PMM) framework helps mitigate the impact of some erroneous predictions by favoring sensor measurements, a single, highly confident but incorrect prediction can still adversely affect the map’s quality and the robot’s subsequent planning.
Finally, while our approach improved map quality and traversability, it did not always result in a complete exploration of the environment, as seen in the case of the SS-frontier approach (Figure 13). This suggests a potential mismatch between the current frontier detection algorithms, which may be confused by the predicted occupancy values, and the generative map. Future work addressing frontier exploration methods for generative occupancy could ensure that the robot can successfully complete exploration.
@@ 10 Conclusion
2020404060608080100100101020203030Denoising StepsFIDKID ($\times 1000$)
2244668810101010151520202525Guidance ScaleFIDKID ($\times 1000$)
Cond = 1 Inpaint = 0 Cond = 0 Inpaint = 1 Cond = 1 Inpaint = 1 0202040406060Values
@@
This ablation study was first presented in [reed2024scenesense] where the SceneSense occupancy prediction network was a conditional diffusion model (as opposed to the unconditional model presented here). When examining these results it was concluded that in this particular application the conditional network does not provided substantial benefits over the equivalent unconditional network.
@@
The number of diffusion steps defines the size of each diffusion step during the reverse diffusion process. Generating reasonable results using the fewest possible denoising steps is desirable behavior to reduce computation time. Additionally too many denoising steps have been shown to introduce sampling drift which results in decreased performance [chen2022diffusiondet, ji2023ddp]. As shown in 15 (a) our method saw the best results when configured to 30 denoising steps. Too few denoising steps results in the network being unable make accurate predictions over the large time step. Increasing the number of denoising steps keeps results relatively stable over time, however you can see the KID is slightly worse using more steps due to sampling drift. Sampling drift is a result of the discrepancy between the distribution of the training and the inference data. During training, the model is trained to reduce a noisy map $x_{t}$ to a ground truth map $x$, at inference time the model iteratively removes noise from its already imperfect noise predictions. These predictions will drift away from the initial corruption distribution which becomes more pronounced at smaller time steps due to the compounding error.
@@
The guidance scale $s$ as defined in [ho2020denoising] is a constant that multiplies the difference of the conditional diffusion and unconditional diffusion to “push” the diffusion process towards the conditioned answer. Setting $s$ too high results in too large of pushes away from reasonable predictions and results in poor generalization to new environments. The best FID is measured when $s=3$ however KID is slightly lower when $s=1$. Further, when examining chart (c) it is shown that the results when conditioning is removed all together ($s=0$ ) are very similar to the best results seen with conditioning enabled (albeit slightly worse). This is likely because most of the useful conditioning information is captured in the local occupancy data, and mapping measured RGB-D points from areas in front of the geometry to under or behind the platform is a very difficult task. The performance of the conditioning only data may be seen to have a larger impact on the overall results if the sensor could capture more local information, such as given a wider FOV or different mounting angle.
@@
This ablation study was first presented in [reed2024online] to evaluate the inference time gains of removing cross-attention based conditioning from the SceneSense occupancy prediction network.
In this ablation we implement the same denoising network structure presented in [reed2024scenesense]. It is a U-net constructed from the HuggingFace Diffusers library of blocks [von-platen-etal-2022-diffusers] and consists of Resnet [he2015resnet] downsampling/upsampling blocks. The diffusion model is trained using randomly shuffled pairs of ground truth local occupancy maps $x$. We use Chameleon cloud computing resources [keahey2020chameleon] to train our model on one A100 with a batch size of $32$ for $250$ epochs or $88,208$ training steps. We use a cosine learning rate scheduler with a 500 step warm up from $10^{-6}$ to $10^{-4}$. The noise scheduler for diffusion is set to $1,000$ noise steps.
@@
We evaluate the inference time of the unconditional diffusion model against the inference time of the conditional model presented in the original SceneSense paper [reed2024scenesense]. The cross-attention enabling trainable parameters are removed for the unconditional model, but the number of output channels for the constructed U-net are held constant between both models. As the ablation results of the original paper show minor, or no performance gain between the conditional and unconditional model in this configuration we do not evaluate the results of the model predictions in these experiments.
@@
Cond. Model [reed2024scenesense] Uncond. Model